Bramble Company has an old factory machine that cost $59,750. The machine has accumulated depreciation of $33,460. Bramble has decided to sell the machine. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

(a) What entry would Bramble make to record the sale of the machine for $29,875 cash?
(b)

What entry would Bramble make to record the sale of the machine for $17,925 cash?No. Account Titles and Explanation Debit Credit (a) Cash Accumulated Depreciation Equipment (b)
Answer #1

a.

Cost of machine = $59,750

Accumulated depreciation = $33,460

Book value of machine = Cost of machine - Accumulated depreciation

= 59,750-33,460

= $26,290

Sale price of machine = $29,875

Gain on sale of machine = Sale price of machine - Book value of machine

= 29,875-26,290

= $3,585

Journal

General Journal Debit Credit
Cash $29,875
Accumulated depreciation - equipment $33,460
Equipment $59,750
Gain on sale of equipment $3,585

b.

Sale price of machine = $17,925

Loss on sale of machine = Book value of machine - Sale price of machine

= 26,290-17,925

= $8,365

Journal

General Journal Debit Credit
Cash $17,925
Accumulated depreciation- equipment $33,460
Loss on sale of equipment $8,365
Equipment $59,750

Kindly comment if you need further assistance. Thanks‼!